Question 6

In a (very fertile) parakeet, brave (B) is dominant to spooky (b); and happy (H) is dominant to sad (h). A dihybrid (B/b·H/h) parakeet was created by mating two true-breeding strains.

When the dihybrid was test-crossed, the following 200 progeny were obtained:

9 brave and happy,

90 brave and sad,

94 spooky and happy,

7 spooky and sad.

c) What is the measured recombination frequency (%) between the B and H loci?

Question 6 options:

49.5

92

51

50.5

8

7

9
Answer #1

Recombination frequency = # recombinants / # total progenies

% RF = RF × 100

# recombinant progenies (minimum in #) = 9 + 7 = 16

# total progenies = 90 + 94 + 9 + 7 = 200

% RF = 16 / 200 × 100 = 8%

ANSWER IS 8.

Recombination the process of exchange of genetic material between non sister chromatids of homologous chromosomes. It occurs during prophase 1 of meiosis 1. It results in the formation of recombinant gametes. Because the chances of recombination is very less (depends upon the distance between the genes undergoing recombination), therefore the number of recombinant progenies is also very less.

Please rate high.